OpenStudy (anonymous):

Please help me! Is (-2, -6) a solution to the equation y=3x yes or no?

jimthompson5910 (jim_thompson5910):

(-2,-6) is an ordered pair in the form (x,y) so x = -2 and y = -6

OpenStudy (anonymous):

Yes \[y=3x\]\[-6=3 \times -2\]\[-6=-6\]

jimthompson5910 (jim_thompson5910):

Plug those values into the equation to get... y = 3x y = 3*(-2) ... replace x with -2 (since x = -2) -6 = 3*(-2) ... replace y with -6 (since y = -6) -6 = -6 ... multiply on the right side Since that last equation is true, this means x = -2 and y = -6 make y = 3x true So (-2,-6) is indeed a solution (one of infinitely many)

OpenStudy (anonymous):

Thank you all so much!

jimthompson5910 (jim_thompson5910):

you're welcome